OpenStudy (anonymous):

Write an equation of a line passing through the points (-2, -1) and (0, 4). Write the final answer in slope-intercept form.

OpenStudy (nirmalnema):

4+1/0+2 =5/2=slope y+1=5/2(x+2) 2y+2=5x+10 5x-2y+8=0

OpenStudy (nirmalnema):

formula for finding slope =(y2-y1)/(x2-x1)

OpenStudy (nirmalnema):

@spencerbeymer

OpenStudy (nirmalnema):

do u understand/...!!

OpenStudy (anonymous):

i know the correct answer

OpenStudy (anonymous):

i wil tell the answer when somebody asks me to

OpenStudy (anonymous):

whats the answer @theGreatStupidPerson

OpenStudy (anonymous):

@nirmalnema the answer is....5x-2y+8=0 correct?

OpenStudy (nirmalnema):

yup..